OpenStudy (anonymous):

Find the Mean, Median, Mode, and Range of the following data: 6, 9, 3, 18, 25, 7, 6, 11, 5

OpenStudy (kirbykirby):

Mean: is the avg of your data: \[ \large \bar{x}=\frac{1}{n}\sum_{i=1}^nx_i\] where \(x_i\) is each data point, and \(n\) is the number of data points you have (here you have 9 data points, so \(n=9\). Mode: is the most frequently occurring data point (which one appears the most often) Range: is the difference between the largest value in your data minus the smallest value range = \(\large x_{\text{largest}} -x_{\text{smallest}}\)

OpenStudy (kirbykirby):

Oops I forgot the median.. the median is the middle of your data once you order it from smallest to largest. So your data set when re-ordered is: 3, 5, 6, 6, 7, 9, 11, 18, 25 If the number of data points is odd, take the middle value from the set If the number of data points is even, take the TWO middle points, and take their mean.

OpenStudy (anonymous):

the median is 25 mode is 6 mean is 10 range is 25,18,11,9,6,5,3 those are your answers :)

OpenStudy (kirbykirby):

the median is certainly not 25

OpenStudy (kirbykirby):

When you find the median, you must sort your data from smallest to largest. Otherwise the median of a data set would not be unique

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

median is in tha middle im not stupid!

OpenStudy (kirbykirby):

but it's the middle in a SORTED list

OpenStudy (kirbykirby):

Even here they tell you so: http://www.mathsisfun.com/definitions/median.html

OpenStudy (anonymous):

DUHHHHH ITS THE MIDDLE NUMBER DUDE

OpenStudy (anonymous):

u callin me a liar?

OpenStudy (kirbykirby):

Same thing with the range.. you don't take the last - first value... you take the LARGEST - SMALLEST value

OpenStudy (anonymous):

guys calm down

OpenStudy (kirbykirby):

It's the middle number when you SORRTTTTTTTTTT the data set

OpenStudy (anonymous):

THATS THE MEAN

OpenStudy (anonymous):

NOT THE MEDIAN

OpenStudy (kirbykirby):

No it is not!! Where on Earth did you learn this?

OpenStudy (anonymous):

fine go with what u think ok? goodbye bro

OpenStudy (kirbykirby):

Please go read this: http://www.purplemath.com/modules/meanmode.htm

OpenStudy (kirbykirby):

It's not what I think, it's fact.

OpenStudy (anonymous):

so what Happie is not right?

OpenStudy (kirbykirby):

Yes unfortunately they are incorrect. You can google the median all you want, and you see that you must re-order the data from smallest to largest, and then find the middle value in that sorted list.

OpenStudy (anonymous):

just go with what u kirby if u want then just... im done

OpenStudy (anonymous):

k kirby than whats the right answer

OpenStudy (kirbykirby):

\[\text{mean}=\bar{x}=\frac{1}{9}(6+9+ 3+ 18+ 25+ 7+ 6+ 11+ 5)\] mode = 6 median: re-sort your data to give: 3, 5, 6, 6, 7, 9, 11, 18, 25... and take the middle value 7 range: largest - smallest value = 25 - 3 = 22

OpenStudy (anonymous):

Mason owns a collection of hardback books in a series about American history. Each book has dimensions as shown below. There are 7 books in the series, stacked side by side. What is the volume of the entire collection? a) one hundred sixty-one cubic inches b) eight hundred forty cubic inches c) nine hundred twenty-four cubic inches d) nine hundred sixty cubic inches

OpenStudy (anonymous):

ill help u the right way

OpenStudy (anonymous):

3,5,6,6,7,9,11,18,25 mean: 10 mode: 6 range: 22 medium: 7